### v2

# Magnon Blockade with Skyrmion Qubit-Magnon Coupling in a Hybrid Quantum System

Abstract
Magnon blockade is a fundamental quantum phenomenon for generating single-magnon state, which gradually becomes one of the candidates for quantum information processing. In this paper, we propose a theoretical scheme to generate the magnon blockade in a hybrid system consisting of a YIG micromagnet and a skyrmion. Considering weak probing of the magnon and driving of the skyrmion qubit, the second-order correlation function is analytically derived, and the optimal condition for realizing the magnon blockade is identified. Under the optimal condition, we systematically analyze the behavior of the second-order correlation function $g^{(2)}(0)$ under different parameter regimes. Our analysis shows that with appropriate driving and probing field intensities, the magnon blockade effect can be significantly enhanced, effectively suppressing multi-magnon states and facilitating the generation of high-purity single-magnon states exhibiting pronounced antibunching. Furthermore, we explore the physical mechanisms underlying the magnon blockade, revealing the coexistence and interplay of conventional and unconventional magnon blockade. This scheme provides a versatile all-magnetic platform for generating high purity single-magnon sources.
I Introduction
Hybrid quantum systems combine the advantages of different subsystems, playing a key role in quantum information processing, networking, and sensing Wallquist et al. (2009); Kimble (2008); Xiang et al. (2013); Kurizki et al. (2015). Examples of hybrid systems are magnonic systems Chen et al. (2023); Xiong et al. (2023); Zuo et al. (2024); Osada et al. (2016); Wolski et al. (2020); Trifunovic et al. (2013); Fukami et al. (2021), optomechanical systems Kippenberg and Vahala (2008); Meystre (2013); Aspelmeyer et al. (2014); Bowen and Milburn (2015); Lai et al. (2022), and photonic-qubit systems Zeng et al. (2023), and so on. Magnons excited in YIG materials, with high spin density, frequency tunability, long lifetime, and enhanced coherence, are promising quantum information carriers Serga et al. (2010); Chumak et al. (2015); Zheng et al. (2023); Yuan et al. (2022); Rameshti et al. (2022); Huebl et al. (2013); Tabuchi et al. (2014); Li et al. (2019); Wang et al. (2016); Zhang et al. (2022). The unique properties of magnons have facilitated hybrid magnonic systems like cavity optomagnonic systems Chen et al. (2023); Xiong et al. (2023); Zuo et al. (2024), cavity magnonic systems Osada et al. (2016); Zhang et al. (2023) and qubit-magnon systems Wolski et al. (2020); Trifunovic et al. (2013); Fukami et al. (2021). In these systems, magnon modes effectively couple with quantum information carriers, including phonon modes, microwave photons, optical photons, and superconducting qubits Osada et al. (2018); Zhang et al. (2016); Shen et al. (2022); Mergenthaler et al. (2017); Parvini et al. (2020); Kounalakis et al. (2022); Liu et al. (2019); Xie et al. (2020); Xu et al. (2023). Nontrivial quantum phenomena in these configurations highlight the potential of hybrid magnonic systems as platforms for exploring and controlling complex quantum interactions He et al. (2023a, 2024); Zhang et al. (2024a); Xu et al. (2023); Liu et al. (2019); Xie et al. (2020); Lu et al. (2021); Xiong (2023); Mukhopadhyay et al. (2022); Liu et al. (2023); Ren et al. (2022); Wang et al. (2019); Zhao et al. (2020a); Hei et al. (2021); Yuan et al. (2020); Qi and Jing (2022); Hisatomi et al. (2016); Zhao et al. (2020b); jun Xu et al. (2021); Wu et al. (2021); Wang et al. (2022); Jin and Jing (2023); Yuan and Duine (2020); Yan et al. (2024); Jin and Jing (2024); Wang et al. (2020); Zhu et al. (2021); Lu et al. (2023); Wu et al. (2024); Couteau et al. (2023); Wang et al. (2024); He et al. (2023b).
Skyrmions, as magnetic nanostructures with particle-like characteristics, exhibit novel magnetic effects Reichhardt et al. (2022); Everschor-Sitte et al. (2018); Liensberger et al. (2021); Hirosawa et al. (2022). Their most notable attribute is their topological protection Ochoa and Tserkovnyak (2019); Back et al. (2020), which enables skyrmions to withstand external disturbances. This topological protection enables state manipulation with ultra-low currents and modulation of electric and magnetic fields Psaroudaki and Panagopoulos (2021); Pan et al. (2024a); Hsu et al. (2016); Birch et al. (2024), facilitating logical transitions within a richly anharmonic operational regime and making them robust information carriers Lonsky and Hoffmann (2020); Reichhardt et al. (2022). Consequently, skyrmions hold significant promise for applications in spintronics and quantum computing Bogdanov and Hubert (1994); Rößler et al. (2006); Mühlbauer et al. (2009); Yu et al. (2010); Jonietz et al. (2010); Schulz et al. (2012); Huang and Chien (2012); Nagaosa and Tokura (2013); Fert et al. (2013, 2017); Bhowal and Spaldin (2022). Experimentally, skyrmions can be generated in centrosymmetric crystals, such as Gd2PdSi3 Kurumaji et al. (2019). In frustrated magnets, skyrmions exhibit internal degrees of freedom related to helicity Leonov and Mostovoy (2015); Lin and Hayami (2016); Kurumaji et al. (2019); Zhang et al. (2017); Yao et al. (2020); Xia et al. (2023). By quantizing collective helicity coordinates, two types of qubits can be constructed Psaroudaki and Panagopoulos (2021), with skyrmion qubits achieving microsecond-scale coherence times, making them suitable for quantum state manipulation. Tunable coherent and dissipative coupling has been achieved between YIG magnon modes and the quantized helicity of skyrmions in skyrmion qubit-magnon systems Pan et al. (2024a). This robust all-magnetic coupling mediates long-range coherent interactions between a magnon and skyrmion qubits, thereby realizing the scalability of skyrmion qubits and providing a promising platform for exploring quantum effects.
In particular, magnon blockade (MB) has attracted special attention as a purely quantum effect Wang et al. (2024); Xu et al. (2023); Liu et al. (2019); Xie et al. (2020); Zhao et al. (2020b); jun Xu et al. (2021); Wu et al. (2021); Wang et al. (2022); Jin and Jing (2023); Yuan and Duine (2020); Yan et al. (2024); Jin and Jing (2024). The MB effect arises from two mechanisms: energy-level anharmonicity, which inhibits the excitation of a second magnon after the first magnon is excited to the single-magnon state, termed conventional magnon blockade (CMB) Liu et al. (2019); Xie et al. (2020); Jin and Jing (2023); Yan et al. (2024), and destructive interference between transition pathways to the two-magnon state, known as as unconventional magnon blockade (UMB) Xie et al. (2020). The MB effect has been theoretically studied in hybrid ferromagnet-superconductor systems Liu et al. (2019). Subsequent studies proposed schemes to achieve antibunched light via MB, including quantum-interference-enhanced MB in YIG-superconducting circuits Xie et al. (2020), and tunable MB in ferromagnet-superconductor systems with dual qubits Li et al. (2021), and enhanced MB in optomechanical-magnetic systems using quantum destructive interference among three indirect transition paths Zhang et al. (2024b). The standard indicator for verifying a stable single-magnon state is $g^{(2)}(0)\rightarrow 0$ Rabl (2011); Liu et al. (2010); Zheng et al. (2019). Analogous to photon blockade Wang et al. (2020); Zhu et al. (2021); Lu et al. (2023); Wu et al. (2024); Couteau et al. (2023); Wu et al. (2024); Lu et al. (2022); Flayac and Savona (2017); Huang et al. (2018) and phonon blockade Xie et al. (2017); Xu et al. (2016); Zeng et al. (2021), MB enables the preparation of single-magnon states, offering the possibility of quantum control at the single-magnon level.
In this work, we theoretically study the realization of MB in YIG ferromagnetic insulators. The system comprises a YIG micromagnet strongly coupled to a resonant skyrmion qubit Pan et al. (2024a). Probing and driving fields are applied to the YIG and the skyrmion qubit, with the latter modulated by electric fields. Building on this setup, we analyze the steady-state properties of the coupled system and derive the optimal MB condition, $\Delta/\gamma\approx\pm\frac{1}{2}g_{ms}/\gamma$, which is consistent with numerical simulations Liu et al. (2019); Carmichael (1999). Using the optimal condition, we investigate the effect of the probing-to-driving field strength ratio on the MB effect. At a ratio of approximately 6, the second-order correlation function $g^{(2)}(0)\sim 10^{-7}$, indicating the strongest MB effect. This result underscores the critical role of the probing-to-driving field strength ratio in optimizing the MB effect. Furthermore, we investigate the mechanisms of MB, revealing the coexistence of CMB and UMB. CMB arises from the intrinsic anharmonicity and energy level splitting in the magnon-qubit system, where the non-equidistant energy spectrum inhibits multi-magnon excitations. UMB is induced by quantum interference effects between transition pathways under the influence of both driving and probing fields. Together, these mechanisms ensure the suppression of multi-magnon states and enhance the MB effect. In another scenario, the skyrmion qubit is exclusively modulated by magnetic fields, and the tprobing is restricted to magnons. Here, MB persists with $g^{(2)}(0)\sim 10^{-2}$, and the optimal condition $\Delta/\gamma\approx\pm\frac{1}{2}\tilde{g}_{ms}/\gamma$ remains valid. Therefore, our study provides theoretical support for the generation of a single-magnon source, which is essential for investigating fundamental aspects of quantum magnons and for potential applications in magnon-based quantum simulation and hybrid quantum devices Li et al. (2019); Hou and Liu (2019); Pan et al. (2024a).
II Model and HAMILTONIAN
As shown in Fig.  1, we study a hybrid quantum system, integrating magnon modes in a YIG micromagnet and skyrmion modes beneath it. Strong coupling between the skyrmion qubit and magnon can be tuned by varying their vertical separation and the radius of the YIG sphere, as the coupling strength depends on both factors Pan et al. (2024a). Within the YIG sphere, spin wave modes with long coherence times can be induced through the application of a uniform bias magnetic field $B_{m}$ Pan et al. (2024a); Gonzalez-Ballestero et al. (2020a, b). The magnon represents spin waves, collective excitations of spin ensembles in magnetic materials, while the skyrmion is a topological excitation characterized by rotations or twists in the spin structure. The coupling between the magnon and the skyrmion arises from the interaction between the magnon’s collective spin wave oscillation and the skyrmion’s localized spin motion within its topological structure.
Figure 1: Schematic representation of the hybrid magnon-Skyrmion system.The magnon is probed by an external microwave field $\Omega_{m}$, while the skyrmion is driven by a microwave field $\Omega_{s}$. The magnon mode interacts with the skyrmion mode through a coupling strength $g_{ms}$. The dissipative processes for both the magnon and skyrmion modes are represented by the damping rates $\kappa_{m}$ and $\kappa_{s}$, respectively.
Here, we focus on the Kittel mode, which features the coherent precession of all spins in the micromagnet, with spins maintaining identical phase and amplitude. The free Hamiltonian of the magnon mode is $\hat{H}_{m}=\omega_{m}\hat{m}^{\dagger}\hat{m}$ (with $\hbar=1$), where $\hat{m}^{\dagger}$ and $\hat{m}$ are the creation and annihilation operators of the magnon mode, respectively, and $\omega_{m}=\gamma_{e}B_{m}$ is the resonance frequency, with $\gamma_{e}$ being the gyromagnetic ratio Pan et al. (2024a). By quantizing the collective coordinates, the skyrmion can be treated as a qubit that can be controlled and manipulated. The energy levels of the $\hat{S}_{z}$ skyrmion qubit are anharmonic Psaroudaki and Panagopoulos (2021); Pan et al. (2024a); Psaroudaki et al. (2023); Pan et al. (2024b), the effective Hamiltonian of the $\hat{S}_{z}$ qubit within the subspace $\{|0\rangle,|1\rangle\}$ is $\hat{H}_{S_{z}}=\frac{K_{0}}{2}\hat{\sigma}_{z}-\frac{E_{z}}{2}\hat{\sigma}_{x}$. The parameters $K_{0}$ and $E_{z}$ are externally modulated, with $K_{0}$ controlled by magnetic fields and $E_{z}$ by electric fields, enabling the manipulation of the qubit’s states Psaroudaki and Panagopoulos (2021). The interaction Hamiltonian is $\hat{H}_{\text{int}}=\frac{\tilde{g}_{ms}}{2}(\hat{m}+\hat{m}^{\dagger})\hat{\sigma}_{x}+\frac{g_{ms}}{2}(\hat{m}+\hat{m}^{\dagger})\hat{\sigma}_{z}$, where $\tilde{g}_{ms}$ describes the coupling between the X- and Y-direction magnetic field generated by the YIG sphere and the skyrmion qubit, and $g_{ms}$ describes the coupling between the Z-direction magnetic field from the YIG sphere and the skyrmion qubit. The maximum coupling strength of $\tilde{g}_{ms}$ can reach up to 40 MHz, whereas $g_{ms}$ can achieve up to 200 MHz Pan et al. (2024a).
To achieve and dynamically manipulate MB, we introduce a microwave drive applied to the skyrmion qubit with a driving frequency $\omega_{s}$. The corresponding Hamiltonian is $H_{s}=\Omega_{s}\cos(\omega_{s}t)\hat{\sigma}_{z}$ Pan et al. (2024a). Additionally, to observe the MB effect, a weak probe field is introduced into the system, with Hamiltonian $H_{d}=\Omega_{d}(\hat{m}^{\dagger}e^{-i\omega_{d}t}+\hat{m}e^{i\omega_{d}t})$ Fujita et al. (2016); Liu et al. (2019), where $\omega_{d}$ and $\Omega_{d}$ are the probe field’s frequency and intensity, respectively. The total Hamiltonian is given by
$\hat{H}_{\text{total}}=$ $\,w_{m}\hat{m}^{\dagger}\hat{m}+\frac{K_{0}}{2}\hat{\sigma}_{z}-\frac{E_{z}}{2}\hat{\sigma}_{x}$
$+\frac{\tilde{g}_{ms}}{2}(\hat{m}+\hat{m}^{\dagger})\hat{\sigma}_{x}+\frac{g_{ms}}{2}(\hat{m}+\hat{m}^{\dagger})\hat{\sigma}_{z}$
$+\Omega_{s}\cos(\omega_{s}t)\hat{\sigma}_{z}+\Omega_{d}(\hat{m}^{\dagger}e^{-i\omega_{d}t}+\hat{m}e^{i\omega_{d}t}).$ (1)
When the skyrmion qubit operates near the degeneracy point, its Hamiltonian $\hat{H}_{S_{z}}$ can be expressed in terms of the eigenstates $\lvert\psi_{2}\rangle=\frac{\sqrt{2}}{2}\lvert 1\rangle-\frac{\sqrt{2}}{2}\lvert 0\rangle$ and $\lvert\psi_{1}\rangle=\frac{\sqrt{2}}{2}\lvert 1\rangle+\frac{\sqrt{2}}{2}\lvert 0\rangle$. In the subspace spanned by $\{\lvert\psi_{2}\rangle,\lvert\psi_{1}\rangle\}$, the Pauli operators are defined as $\tilde{\sigma}_{z}=\lvert\psi_{2}\rangle\langle\psi_{2}\rvert-\lvert\psi_{1}\rangle\langle\psi_{1}\rvert$ and $\tilde{\sigma}_{x}=\lvert\psi_{2}\rangle\langle\psi_{1}\rvert+\lvert\psi_{1}\rangle\langle\psi_{2}\rvert$. The coupling term $\tilde{g}_{ms}$ does not induce energy level transitions but causes a small shift ($\tilde{g}_{ms}\approx 50.1\,\text{MHz}$) in the skyrmion qubit energy levels. Because this shift is much smaller than the gigahertz-scale resonant frequencies ($E_{z}\approx\omega_{m}\approx 9.8\,\text{GHz}$), it can be neglected Pan et al. (2024a, b). Using the rotating-wave approximation (RWA) and considering the time-evolution operator in the rotating frame, $U=e^{-i\left(\frac{\omega_{s}\tilde{\sigma}_{z}t}{2}+w_{d}t\hat{m}^{\dagger}\hat{m}\right)}$, the hybrid quantum system Hamiltonian can be simplified. Thus, the effective Hamiltonian can be written as
$H_{\text{eff}}=$ $\,\Delta_{m}\hat{m}^{\dagger}\hat{m}+\frac{\Delta_{s}}{2}\tilde{\sigma}_{z}+\frac{g_{ms}}{2}(\hat{m}\tilde{\sigma}_{+}+\hat{m}^{\dagger}\tilde{\sigma}_{-})$
$+\frac{\Omega_{s}}{2}(\tilde{\sigma}_{+}+\tilde{\sigma}_{-})+\Omega_{d}(\hat{m}^{\dagger}+\hat{m}).$ (2)
where $\Delta_{m}\equiv\omega_{m}-\omega_{d}$ ($\Delta_{s}\equiv E_{z}-\omega_{s}$) represents the detuning between the magnon mode and the probe field frequency, while $\Delta_{s}$ represents the detuning between the skyrmion qubit frequency and the driving frequency. We choose the parameter conditions as $\omega_{d}\approx\omega_{s}$, $\omega_{d}\gg g_{ms}$, and $\omega_{s}\gg\Omega_{s}$, with $\Delta=\Delta_{s}=\Delta_{m}$, which can be flexibly tuned in the current experimental setup.
III Magnon Blockade and OPTIMAL CONDITIONS
III.1 The solution of the second-order correlation function
In the weak-driving regime, where $\Omega_{d},\Omega_{s}\ll k_{m},k_{s}$, and the effects of quantum jumps on short time scales are neglected, the wave function of the composite system can be approximated within the framework of an analytical model Wu et al. (2021); Liu et al. (2019); Xie et al. (2020); Wang et al. (2022); Kómár et al. (2013); Flayac and Savona (2017); Tang et al. (2019); Jin and Jing (2024, 2023). In particular, the dynamics of the system can be described within an approximate framework using a non-Hermitian Hamiltonian
$\hat{H}_{\mathrm{non}}=$ $\,\hat{H}_{\mathrm{eff}}-i\frac{k_{m}}{2}m^{\dagger}m$ $-i\frac{k_{s}}{2}\tilde{\sigma}_{+}\tilde{\sigma}_{-}$ (3)
thus, the Hilbert space of the composite system can be truncated to a subspace corresponding to a few excitations, and the system’s evolution can be effectively approximated by a pure state, as described below
$|\psi\rangle=$ $\,C_{g^{\prime},0}|g^{\prime},0\rangle+C_{g^{\prime},1}|g^{\prime},1\rangle+C_{e^{\prime},0}|e^{\prime},0\rangle$
$+C_{e^{\prime},1}|e^{\prime},1\rangle+C_{g^{\prime},2}|g^{\prime},2\rangle$ (4)
where $C_{m^{\prime},s}$ ($m=g^{\prime},e^{\prime}\;\text{and}\;s=0,1,2$) are the probability amplitudes for the states $|ms\rangle$, and $|g(e),n\rangle$ represents the ground (excited) state of the skyrmion qubit and the Fock state of the magnon mode, respectively.
Based on the Schrödinger equation $i\frac{d}{dt}|\psi(t)=\hat{H}_{\text{non}}|\psi(t)\rangle$, the dynamical evolution of the coefficients $C_{ms}$ can be derived from the equations. The corresponding dynamical equations can then be written as
$i\dot{C}_{g^{\prime},0}$ $=\frac{\Omega_{s}}{2}C_{e^{\prime},0}+\Omega_{d}C_{g^{\prime},1}$ (5a)
$\vspace{-1em}i\dot{C}_{e^{\prime},0}$ $=(\Delta_{s}-i\frac{k_{s}}{2})C_{e^{\prime},0}+\frac{g_{ms}}{2}C_{g^{\prime},1}+\frac{\Omega_{s}}{2}C_{g^{\prime},0}$ (5b)
$\vspace{-1em}i\dot{C}_{g^{\prime},1}$ $=(\Delta_{m}-i\frac{k_{m}}{2})C_{g^{\prime},1}+\frac{g_{ms}}{2}C_{e^{\prime},0}+\Omega_{d}C_{g^{\prime},0}$ (5c)
$\vspace{-1em}i\dot{C}_{e^{\prime},1}$ $=(\Delta_{m}+\Delta_{s}-i\frac{k_{m}}{2}-i\frac{k_{s}}{2})C_{e^{\prime},1}+\frac{\sqrt{2}}{2}g_{ms}C_{g^{\prime},2}$
$\quad+\frac{\Omega_{s}}{2}C_{g^{\prime},1}+\Omega_{d}C_{e^{\prime},0}$ (5d)
$\vspace{-1em}i\dot{C}_{g^{\prime},2}$ $=(\Delta_{m}-i\frac{k_{m}}{2})C_{g^{\prime},2}+\sqrt{2}\Omega_{d}C_{g^{\prime},1}+\frac{\sqrt{2}}{2}g_{ms}C_{e^{\prime},1}\vspace{-1em}$ (5e)
then its steady-state solution can be obtained from
00 0 $=\frac{\Omega_{s}}{2}C_{e^{\prime},0}+\Omega_{d}C_{g^{\prime},1},$ (6a)
00 0 $=(\Delta_{s}-i\frac{k_{s}}{2})C_{e^{\prime},0}+\frac{g_{ms}}{2}C_{g^{\prime},1}+\frac{\Omega_{s}}{2}C_{g^{\prime},0},$ (6b)
00 0 $=(\Delta_{m}-i\frac{k_{m}}{2})C_{g^{\prime},1}+\frac{g_{ms}}{2}C_{e^{\prime},0}+\Omega_{d}C_{g^{\prime},0},$ (6c)
00 0 $=(\Delta_{m}+\Delta_{s}-i\frac{k_{m}}{2}-i\frac{k_{s}}{2})C_{e^{\prime},1}+\frac{\sqrt{2}}{2}g_{ms}C_{g^{\prime},2}$
$\quad+\frac{\Omega_{s}}{2}C_{g^{\prime},1}+\Omega_{d}C_{e^{\prime},0},$ (6d)
00 0 $=(\Delta_{m}-i\frac{k_{m}}{2})C_{g^{\prime},2}+\sqrt{2}\Omega_{d}C_{g^{\prime},1}+\frac{\sqrt{2}}{2}g_{ms}C_{e^{\prime},1}.$ (6e)
In the weak-driving regime ($\Omega_{d},\Omega_{s}\ll\kappa_{m},\kappa_{s}$), where $|C_{g^{\prime},0}|\gg|C_{g^{\prime},1}|,|C_{g^{\prime},2}|$ and $|C_{e^{\prime},0}|\gg|C_{g^{\prime},3}|,|C_{e^{\prime},1}|,|C_{e^{\prime},2}|$. By applying perturbation theory, we neglect the higher-order terms that contribute minimally in the equation of motion for the lower-order variables. We can derive the steady-state solutions for the probability amplitudes
$C_{g^{\prime},0}$ $\approx 1$ (7a)
$C_{e^{\prime},0}$ $=\frac{-2\Omega_{d}g_{ms}+2\tilde{\Delta}_{m}\Omega_{s}}{-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2}}$ (7b)
$C_{g^{\prime},1}$ $=\frac{4\tilde{\Delta}_{s}\Omega_{d}-\Omega_{s}g_{ms}}{-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2}}$ (7c)
$C_{e^{\prime},1}$ $=\frac{-\Omega_{s}DA-2A\Omega_{d}E+8\tilde{\Delta}_{s}\Omega_{d}g_{ms}-2g_{ms}^{2}\Omega_{d}\Omega_{s}}{(-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2})(2AB-g_{ms}^{2})}$ (7d)
$C_{g^{\prime},2}$ $=\frac{-2\sqrt{2}BC+\Omega_{s}Dg_{ms}+2\Omega_{d}Eg_{ms}}{(-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2})(2\sqrt{2}AB-\sqrt{2}g_{ms}^{2})}$ (7e)
where $\tilde{\Delta}_{m}=\Delta_{m}-i\frac{k_{m}}{2},\,\tilde{\Delta}_{s}=\Delta_{s}-i\frac{k_{s}}{2},\,A=2\big{(}\Delta_{m}-i\frac{k_{m}}{2}\big{)},\,B=\big{(}\Delta_{m}+\Delta_{s}\big{)}-i\frac{k_{m}}{2}-i\frac{k_{s}}{2},\,C=4\sqrt{2}\tilde{\Delta}_{s}\Omega_{d}^{2}-\sqrt{2}g_{ms}\Omega_{d}\Omega_{s},\,D=4\tilde{\Delta}_{s}\Omega_{d}-\Omega_{s}g_{ms},\,E=-2\Omega_{d}g_{ms}+2\tilde{\Delta}_{m}\Omega_{s}.$
With the steady-state amplitudes in Eqs. (7c) and (7e), the function $g^{(2)}$ can then be expressed by
$g^{(2)}(0)$ $\approx\frac{2\left|C_{g^{\prime},2}\right|^{2}}{\left|C_{g^{\prime},1}\right|^{4}}=\frac{2\left|\eta_{1}\right|^{2}\cdot\left|\chi\right|^{2}}{\left|\eta_{2}\right|^{2}\left|D\right|^{4}}$ (8)
in which $\eta_{1}=-2\sqrt{2}BC+\Omega_{s}Dg_{ms}+2\Omega_{d}Eg_{ms},\eta_{2}=2\sqrt{2}AB-\sqrt{2}g_{ms}^{2},\chi=-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2}$.
The population in the state $|g^{\prime},1\rangle$ can be optimized by
$\Delta_{s}\Delta_{m}$ $\approx\tilde{\Delta}_{s}\tilde{\Delta}_{m}=\frac{1}{4}g_{ms}^{2}$ (9)
III.2 Mechanisms of The Magnon Blockade
Figure 2: (a) Energy spectrum of the skyrmion qubit-magnon system, showing vacuum Rabi splitting induced by magnon-qubit coupling. (b) Transition pathways in the skyrmion qubit-YIG system.
To explore the mechanism of magnon blockade in a skyrmion-qubit-magnon system, we compute the eigenfunctions and examine the energy spectrum. The ground state is $|g^{\prime},0\rangle$ with eigenvalue $E_{0}=0$, while for $n\geq 1$, the excited (dressed) states are expressed as
$|\psi_{n,\pm}\rangle=C_{g^{\prime},n}|g^{\prime},n\rangle+C_{e^{\prime},n-1}|e^{\prime},n-1\rangle,$
where the normalized coefficients are
$C_{g^{\prime},n}=\pm\frac{g_{ms}\sqrt{n}}{\sqrt{g_{ms}^{2}n+4[n\omega_{m}-E_{n,\pm}]^{2}}},$
and
$C_{e^{\prime},n-1}=\pm\sqrt{\frac{4[n\omega_{m}-E_{n,\pm}]^{2}}{g_{ms}^{2}n+4[n\omega_{m}-E_{n,\pm}]^{2}}}.$
The associated eigenvalues for $\psi_{n,\pm}$ are given by
$E_{n,\pm}=\frac{(2n-1)\omega_{m}+E_{z}}{2}\pm\frac{\sqrt{[\omega_{m}-E_{z}]^{2}+g_{ms}^{2}n}}{2}.$
These coefficients and eigenvalues describe the dressed states in the system and ensure their normalization.
Under the strong-coupling regime, the real parts of the eigenvalues provide an accurate description of the energy spectrum, while the imaginary components, associated with dissipation, can be considered negligible del Valle et al. (2009).
For the case ($\omega_{m}=E_{z}$), the eigenvalues are $E_{n,\pm}=n\omega_{m}\pm\frac{g_{ms}\sqrt{n}}{2}$, and the eigenstates are $|\psi_{n,\pm}\rangle=\frac{1}{\sqrt{2}}|g^{\prime},n\rangle\pm\frac{1}{\sqrt{2}}|e^{\prime},n-1\rangle$. The physical mechanism of the magnon blockade can be explained by the energy-level structure shown in Fig. 2(a). Transitions between the states $|g^{\prime},0\rangle$ and |±$|\pm^{\prime},1\rangle$ occur when the detuning satisfies $\omega_{m}-\omega_{d}=\pm\Delta$ ($\Delta=\frac{1}{2}g_{ms}$). Conversely, transitions from |±$|\pm^{\prime},1\rangle$ to |±$|\pm^{\prime},2\rangle$ are suppressed due to the qubit-induced vacuum Rabi splitting, which requires the condition $2\omega_{m}\pm 2\Delta\neq 2\omega_{m}\pm\frac{\sqrt{2}}{2}g_{ms}$. This suppression arises from the anharmonicity in the energy spectrum caused by the vacuum Rabi splitting. The anharmonicity prevents the second magnon excitation from resonantly reaching the higher states |±$|\pm^{\prime},2\rangle$, confining the system to the single-magnon regime and leading to the CMB effect. In Fig. 2(b), the transition pathways under the combined influence of the probe field and the drive field are depicted. The probe field primarily induces transitions between magnon states, such as $|g^{\prime},0\rangle\xrightarrow{\Omega_{d}}|g^{\prime},1\rangle\xrightarrow{\sqrt{2}\Omega_{d}}|g^{\prime},2\rangle$. Meanwhile, the drive field facilitates transitions between magnon states and qubit states through coupling mechanisms, for example, $|g^{\prime},0\rangle\xrightarrow{\frac{\Omega_{s}}{2}}|e^{\prime},0\rangle\xrightarrow{2^{-1}g_{ms}}|g^{\prime},1\rangle\xrightarrow{\frac{\Omega_{s}}{2}}|e^{\prime},1\rangle\xrightarrow{2^{-1/2}g_{ms}}|g^{\prime},2\rangle$. These pathways may interfere coherently, suppressing the population of the quantum state $|g^{\prime},2\rangle$ and leading to the emergence of the UMB effect.
III.3 Numerical Simulation of the Second-Order Correlation Function
Considering the damping effects induced by the system-bath coupling, the dissipative dynamics of the hybrid skyrmion qubit-magnon system, with a focus on the magnon mode, can be described by the master equation Carmichael (1999)
$\frac{\partial\rho}{\partial t}=$ $-i[H_{\text{eff}},\rho]+\frac{\kappa_{m}}{2}(n_{\text{th}}+1)\mathcal{L}_{m}[\rho]$ (14)
$+\frac{\kappa_{m}}{2}n_{\text{th}}\mathcal{L}^{\prime}_{m}[\rho]+\frac{\kappa_{s}}{2}\mathcal{L}_{\tilde{\sigma}}[\rho],$
Here, $\kappa_{s}$ and $\kappa_{m}$ denote the decay rates of the skyrmion qubit and magnon mode, with $\kappa_{s}=\kappa_{m}$ assumed for simplicity. The Pauli operators $\tilde{\sigma}_{+}=|\psi_{2}\rangle\langle\psi_{1}|$ and $\tilde{\sigma}_{-}=|\psi_{1}\rangle\langle\psi_{2}|$ are defined using the skyrmion qubit eigenstates $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$. The system density matrix is $\rho$, and the Lindblad superoperators are defined below
$\mathcal{L}_{m}[\rho]$ $=2m\rho m^{\dagger}-m^{\dagger}m\rho-\rho m^{\dagger}m,$ (15)
$\mathcal{L}_{\tilde{\sigma}}[\rho]$ $=2\tilde{\sigma}_{-}\rho\tilde{\sigma}_{+}-\tilde{\sigma}_{+}\tilde{\sigma}_{-}\rho-\rho\tilde{\sigma}_{+}\tilde{\sigma}_{-}.$ (16)
The phenomenon of magnon blockade can be quantified using the second-order correlation function, given by Scully and Zubairy (1997)
$g^{(2)}(0)=\frac{\langle m^{\dagger}m^{\dagger}mm\rangle}{\langle m^{\dagger}m\rangle^{2}}.$ (17)
If $g^{(2)}(0)>1$, this indicates that the steady-state magnon number distribution follows super-Poissonian statistics, exhibiting magnon bunching. Conversely, if $10<g^{(2)}(0)<$ 1, the magnon number distribution satisfies sub-Poissonian statistics, known as magnon antibunching Rabl (2011); Liu et al. (2010); Zheng et al. (2019). Specifically, when $g^{(2)}(0)\rightarrow 0$, magnon blockade occurs.
Figure 3: The steady-state equal-time second-order correlation function $\log_{10}g^{(2)}(0)$ varies with the driving detuning $\Delta/\gamma$ and the coupling strength $g_{ms}/\gamma$ in (a), and with the driving field $\Omega_{s}/\gamma$ and the probe field $\Omega_{d}/\gamma$ in (b). Frequencies are normalized to $\gamma=2\pi\times 1$ MHz for consistency. In (a), the parameters are $\Omega_{s}/\gamma=0.06$ and $\Omega_{d}/\gamma=0.01$, while in (b), $\Delta/\gamma=9.8$ and $g_{ms}/\gamma=19.6$. The other parameters are $\kappa_{m}/\gamma=\kappa_{s}/\gamma=1$ and $n_{th}=0$.
To verify the optimal condition in Eq. (9), we numerically solve the master equation (14) by the Python package QuTiP Johansson et al. (2013, 2012). In Fig. 3(a), the steady-state second-order correlation function $\log_{10}g^{(2)}(0)$ depends on the drive detuning $\Delta/\gamma$ and the coupling strength $g_{ms}/\gamma$ between the magnon and the skyrmion qubit. The transition of $g^{(2)}(0)$ from strong quantum correlations to classical correlations is illustrated by the color gradient, which ranges from deep purple to yellow. The system shows pronounced antibunching near the white dashed lines, where $\Delta/\gamma\approx\pm\frac{1}{2}g_{ms}/\gamma$, and $g^{(2)}(0)$ reaches its minimum. This result highlights the emergence of the MB effect, aligning with the analytically derived the optimal condition. Fig. 3(b) provides a detailed analysis of $g^{(2)}(0)$ as a function of the driving field intensity $\Omega_{s}/\gamma$ and the probe field intensity $\Omega_{d}/\gamma$. As $\Omega_{s}/\gamma$ and $\Omega_{d}/\gamma$ increase, $g^{(2)}(0)$ exhibits significant changes and reaches its minimum near the white dashed lines. The results indicate that the ratio $\Omega_{s}/\Omega_{d}\approx 6$ corresponds to the optimal condition for realizing the MB effect.
Figure 4: The steady-state $\log_{10}g^{(2)}(0)$ is plotted as a function of (a) the coupling strength $g_{ms}$ and (b) the driving field intensity $\Omega_{s}/\gamma$, as well as the dissipation rate $\kappa$, under the condition $\Delta/\gamma\approx\frac{1}{2}g_{ms}/\gamma$. In (a), $\Omega_{s}/\gamma=0.06$ and $\Omega_{d}/\gamma=0.01$, while in (b) $g_{ms}/\gamma=19.6$, and $n_{th}=0$.
Both the probe field and the driving field play a crucial role in enhancing the MB effect. In Fig. 4, we investigate the second-order correlation function $g^{(2)}(0)$ as a function of the coupling strength $g_{ms}$, driving field intensity $\Omega_{s}$, and dissipation rate $\kappa$. In Fig. 4(a), $\log_{10}g^{(2)}(0)$ remains extremely low (deep purple) under conditions of high $g_{ms}$ and low $\kappa$, demonstrating a pronounced antibunching effect. With $g_{ms}/\gamma=19.6$ fixed, Fig. 4(b) explores the influence of $\Omega_{s}$ and $\kappa$ on $g^{(2)}(0)$. For $\Omega_{d}/\gamma=0.01$ and $\Omega_{s}/\gamma\approx 0.06$, $\log_{10}g^{(2)}(0)$ achieves its minimum within the dissipation range $\kappa/\gamma$ between 0.1 and 0.6, corresponding to the strongest MB effect. This confirms that $\Omega_{s}/\Omega_{d}\approx 6$ is optimal for $\kappa/\gamma$ in the range 0.1 to 1.2. These results provide the key parameter values for $\Omega_{s}$, $\kappa$, and $\Omega_{d}$ that maximize the MB effect.
Figure 5: The equal-time second-order correlation function $g^{(2)}(0)$ is shown as a function of the normalized detuning $\Delta/\gamma$. As illustrated in (a), for the magnon drive field intensity $\Omega_{d}/\gamma=0.004$, 0.01, and 0.012, the Skyrmion drive field $\Omega_{s}/\gamma$ is kept constant at 0.06. In (b), for the skyrmion drive field intensity $\Omega_{s}/\gamma=0.001$, 0.05, and 0.09, the magnon drive field $\Omega_{d}/\gamma$ is fixed at 0.01. The coupling strength between the magnon and skyrmion modes is $g_{\text{ms}}/\gamma=19.6$. The other system parameters are $\kappa_{m}/\gamma=\kappa_{q}/\gamma=0.15$, and $n_{th}=0$.
Next, we observe a significant dependence of the second-order correlation function $g^{(2)}(0)$ on the driving detuning $\Delta/\gamma$, as shown in in Fig. 5. In Fig. 5(a), the variation of $g^{(2)}(0)$ is shown for different magnon driving field intensities ($\Omega_{d}/\gamma=0.004,0.01,0.012$), with corresponding ratios $\Omega_{s}/\Omega_{d}=15,6,5$, while the skyrmion driving field intensity is fixed at $\Omega_{s}/\gamma=0.06$. Around $\Delta/\gamma\approx 9.8$, $g^{(2)}(0)$ reaches $10^{-7.3}$, indicating a pronounced antibunching effect and a high-purity magnon blockade state. However, around $\Delta/\gamma\approx-9.8$, the antibunching effect becomes less sensitive to modulation, and $g^{(2)}(0)$ reaches $10^{-3.3}$. In Fig. 5(b), $g^{(2)}(0)$ is plotted as a function of the skyrmion driving field intensity ($\Omega_{s}/\gamma=0.001,0.05,0.09$), with corresponding ratios $\Omega_{s}/\Omega_{d}=0.1,5,9$, while the magnon driving field intensity is fixed at $\Omega_{d}/\gamma=0.01$. In this parameter regime, the MB effect is present but weaker, with $g^{(2)}(0)$ reaching $10^{-2.7}$, $10^{-3.2}$, and $10^{-3.4}$, respectively. Notably, in both Fig. 5(a) and Fig. 5(b), the numerical and analytical results agree almost perfectly, demonstrating the reliability of the model and theoretical calculations. Further analysis confirms that the optimal conditions are $\Omega_{s}/\Omega_{d}\approx 6$, $\Delta/\gamma\approx\pm\frac{1}{2}g_{ms}/\gamma$.
Figure 6: The dependence of the second-order correlation function $g^{(2)}(0)$ on both the skyrmion driving field $\Omega_{s}/\gamma$ and the magnon driving field $\Omega_{d}/\gamma$. In (a), $g^{(2)}(0)$ is plotted as a function of $\Omega_{s}/\gamma$ for detuning values $\kappa/\gamma=0.15,0.30,0.45,1$, and $\Omega_{d}/\gamma=0.01$. In (b), the behavior of $g^{(2)}(0)$ is depicted as a function of $\Omega_{d}/\gamma$ under the same detuning conditions, $\Omega_{s}/\gamma=0.06$, while the other parameters are the same as those in Fig. 5.
In Fig. 6, we illustrate the relationship between $\Omega_{d}$, $g_{ms}$, $\Omega_{s}$, and $\kappa$, showing that the system achieves the strongest antibunching effect near $\Omega_{s}/\gamma\approx 0.06$ and $\Omega_{d}/\gamma\approx 0.01$. Stronger antibunching effects are observed at lower dissipation rates, such as $\kappa/\gamma=0.15$, which is evidenced by the significantly reduced values of $g^{(2)}(0)$. As the dissipation rate increases, the antibunching effect weakens; however, $g^{(2)}(0)$ remains at a sufficiently low value even under higher dissipation conditions. For instance, when $\kappa/\gamma=1.00$, $g^{(2)}(0)$ can reach $10^{-4}$.
Our approach to the MB effect aims to drive the magnon system into a quantum state where the first excited level is highly populated, while higher excited levels remain largely unoccupied, thereby suppressing multi-excitation events. This mechanism minimizes the second-order correlation function $g^{(2)}(0)$, a key indicator of the MB effect. It is expressed as $g^{(2)}(0)\approx 2P_{2}/P_{1}^{2}$, which aligns with Eqs. (8) and (17). Fig. 7 shows the population dynamics of the lowest Fock states of the magnon mode under optimal conditions. The population of the single-excitation level $P_{1}$ is approximately $10^{-2}$, while the population of the double-excitation state $P_{2}$ is around $10^{-11}$. With the direct coupling between magnons and qubits, a nearly perfect magnon blockade can be achieved Jin and Jing (2024), characterized by $P_{1}\sim 0.40$ and $g^{(2)}(0)\sim 10^{-5}$. These results provide strong evidence of MB and highlight its potential for precise quantum control in hybrid quantum systems.
Figure 7: The population dynamics of the lowest few Fock states of the magnon mode under the optimal conditions $\Delta/\gamma\approx\pm\frac{1}{2}g_{\text{ms}}/\gamma$ and $\Omega_{s}=6\Omega_{d}$. The parameters are $g_{\text{ms}}/\gamma=19.6$, $\Omega_{d}/\gamma=0.01$, and $\kappa/\gamma=0.15$. The initial state is $\rho(0)=|g^{\prime}0\rangle\langle g^{\prime}0|$.
IV discussions
In the previous analysis, we investigated the case where the skyrmion qubit is solely modulated by an electric field ($E_{z}\approx 9.8\,\text{GHz}$ and $K_{0}\approx 0$) and observed a clear MB effect. This analysis provided valuable insights into the interaction mechanisms under such conditions. Moving forward, we consider another scenario where $E_{z}\approx 0$ and $K_{0}\approx 9.8\,\text{GHz}$. In this configuration, the system is designed to exclusively probe the magnon mode, with the external driving on the skyrmion qubit removed. The Hamiltonian is written as
$\hat{H}_{\text{total}}^{\prime}=$ $\,w_{m}\hat{m}^{\dagger}\hat{m}+\frac{K_{0}}{2}\hat{\sigma}_{z}+\frac{\tilde{g}_{ms}}{2}(\hat{m}+\hat{m}^{\dagger})\hat{\sigma}_{x}$ (18)
$+\frac{g_{ms}}{2}(\hat{m}+\hat{m}^{\dagger})\hat{\sigma}_{z}+\Omega_{d}(\hat{m}^{\dagger}e^{-i\omega_{d}t}+\hat{m}e^{i\omega_{d}t}).$
Following the approach used in the previous analysis for handling the Hamiltonian, The effective Hamiltonian can then be written as
$\hat{H}_{\text{eff}}^{\prime}=$ $\Delta_{m}\hat{m}^{\dagger}\hat{m}+\frac{\Delta_{s}}{2}\sigma_{z}+\frac{\tilde{g}_{ms}}{2}\left(\hat{m}\sigma_{+}+\hat{m}^{\dagger}\sigma_{-}\right)$
$+\Omega_{d}\left(\hat{m}^{\dagger}+\hat{m}\right),$
where $\Delta_{m}\equiv\omega_{m}-\omega_{d}$ ($\Delta_{s}\equiv K_{0}-\omega_{d}$) represents the detuning between the magnon mode and the driving field frequency, while $\Delta_{s}$ represents the detuning between the skyrmion qubit frequency and the driving frequency. Simultaneously, $\Delta=\Delta_{s}=\Delta_{m}$. The statistical properties of the magnon are investigated by numerically solving the master equationCarmichael (1999)
$\frac{\partial\rho}{\partial t}=$ $-i[\hat{H}_{\text{eff}}^{\prime},\rho]+\frac{\kappa_{m}}{2}\mathcal{L}_{m}[\rho]+\frac{\kappa_{s}}{2}\mathcal{L}_{\sigma}[\rho],$ (20)
the dissipator is defined as the Lindblad superoperators $\mathcal{L}[\rho]=2o\rho o^{\dagger}-o^{\dagger}o\rho-\rho o^{\dagger}o$, with $o=\hat{m},\sigma_{-}$ indicating, respectively.
Figure 8: The variation of the steady-state equal-time second-order correlation function $\log_{10}g^{(2)}(0)$ is shown as a function of the driving detuning $\Delta/\gamma$ and the skyrmion qubit-magnon coupling strength $\tilde{g}_{ms}/\gamma$. The probe field intensity of the magnons is fixed at $\Omega_{d}/\gamma=0.01$, while other system parameters are set to $\kappa_{m}/\gamma=\kappa_{q}/\gamma=1$.
The Fig. 8 shows the variation of the steady-state equal-time second-order correlation function $\log_{10}g^{(2)}(0)$ as a function of the driving detuning $\Delta/\gamma$ and the skyrmion qubit-magnon coupling strength $g_{ms}/\gamma$. It can also be observed that under the optimal conditions where $\Delta/\gamma\approx\pm\frac{1}{2}\tilde{g}_{ms}/\gamma$, $\log_{10}g^{(2)}(0)$ reaches -2, indicating the occurrence of a strong MB effect.
Figure 9: The steady-state $\log_{10}g^{(2)}(0)$ is presented as a function of (a) the skyrmion-qubit coupling strength $g_{ms}$ and (b) the control field intensity $\Omega_{s}/\gamma$, along with the dissipation rate $\kappa$, under the condition $\Delta/\gamma\approx\pm\frac{1}{2}\tilde{g}_{ms}/\gamma$. In panel (a), $\Omega_{d}/\gamma=0.01$ is set, while in panel (b), $g_{ms}/\gamma=50.1$ is chosen, and $n_{th}=0$.
In the analysis of Fig. 9(a), the effects of the dissipation rate $k$ and the skyrmion qubit-magnon coupling strength $\tilde{g}_{ms}$ on the system’s behavior are explored. The result indicates that, under high coupling strength and low dissipation rate $k$, $\log_{10}g^{(2)}(0)$ remains at a low value of approximately -5. In Fig. 9(b), the influence of the driving intensity $\Omega_{d}$ on the system is further investigated. When the dissipation rate $k$ is fixed, small variations in the driving intensity (e.g., $0.001\leq\Omega_{d}/\gamma\leq 0.01$) have little effect on the system, with $\log_{10}g^{(2)}(0)$ remaining stable at around -3.5. However, with a larger variation in the driving intensity, the system exhibits pronounced fluctuations, leading to noticeable changes in the second-order correlation function. For instance, as $\Omega_{d}/\gamma$ increases from 0.1 to 0.6, $\log_{10}g^{(2)}(0)$ rises from -3.5 to -1.5. Notably, even in the absence of an external drive applied to the skyrmion qubit, applying a drive solely to the magnons effectively induces a MB effect.
V conclusions
In conclusion, we propose a scheme to realize a high-purity single-magnon state in a skyrmion qubit-magnon system. We consider two scenarios: one where the magnon is probed and the skyrmion qubit is driven, with the skyrmion qubit modulated by an electric field ($E_{z}\approx 9.8\,\text{GHz}$ and $K_{0}\approx 0$); and another where only the magnon is probed, with the skyrmion qubit modulated by a magnetic field ($E_{z}\approx 0$ and $K_{0}\approx 9.8\,\text{GHz}$). In the first scenario, we derive the analytical expression for the second-order correlation function and the optimal conditions through theoretical analysis and numerical simulations. Specifically, the detuning is one-half of the transverse coupling strength of the skyrmion qubit, and the driving field strength applied to the skyrmion qubit is six times the probing field strength applied to the magnon. Under these optimal conditions, and with the thermal magnon occupation number $n_{\text{th}}=0$, a strong MB effect is observed, with the second-order correlation function $g^{(2)}(0)\sim 10^{-7}$. This demonstrates that detuning, driving field strength, and probing field strength significantly influence the enhancement of the MB effect. Moreover, we analyze the physical mechanism responsible for the magnon blockade and confirm the coexistence of CMB and UMB. In the second scenario, by appropriately tuning the coupling strength $g_{\text{ms}}$ and the probing field $\Omega_{d}$, the MB effect is also observed. Further analysis of the system’s dissipation, driving fields shows that, under suitable parameter settings, the MB effect is still observed, with $g^{(2)}(0)$ reaching approximately $10^{-2}$. These findings establish a theoretical framework for studying and advancing the potential realization of single-magnon emitters in hybrid quantum systems.
Additionally, the versatility of this platform allows for further exploration of its applicability in broader quantum technologies. By leveraging the tunability of coupling parameters and system configurations, this scheme may inspire the design of scalable quantum architectures and novel quantum devices. These advancements could contribute to enriching the functionality of hybrid quantum systems and unlocking new possibilities in magnon-based quantum technologies.
Acknowledgments
We thank Feng-Yang Zhang, Mingsong Ding, and Ye-Ting Yan for constructive discussions. This work was supported by the National Natural Science Foundation of China (Grants No. 12274053).